OpenStudy (anonymous):

Let dy/dx=xy . Use Euler's method with step size of 0.5 and the initial condition of y(0) = 3 to estimate y(2)....check my answer? I got 9.84.

OpenStudy (anonymous):

Given the differential equation: dy/dx = y(1+x), y(0)=1, Use Euler's method with step size .1 to approximate y(.3). y' = y(1+x), y'(0) = 1(1+0)=1 ->the solution has slope 1 at the point (0,1); x0=0, y0=1, h=0.1, F(x,y)=y(1+x) y1=y0+h*F(x0,y0) y1=1 + 0.1(1(1+0)) y1=1.01 y2=1...

OpenStudy (anonymous):

I got 9.84 as an answer. Is this right?

OpenStudy (anonymous):

Yes, please give me a medal

OpenStudy (anonymous):

Thank you!